# Market-based insurance ratemaking: application to pet insurance

### v1

@@ 1 Introduction
Our approach assumes that the distribution of the risk $X$ is parameterized by $\theta\in\Theta\subset\mathbb{R}^{d}$. Ideally, if the loading functions $f_{i}$ were known for $i=1,\ldots,n$, a procedure similar to the generalized method of moments could be applied (see Hansen [13]). Unfortunately, these functions are unknown in this context to anyone not internal to the companies from which the premium quotes were obtained. Therefore, in order to proceed we will perform an estimation under a two-stage procedure. We posit a prior distribution $\pi(\theta)$ to delimit the parameter space of the risk model from which we can readily sample $\theta$. We compute the pure premiums $p_{i}^{\theta}=\mathbb{E}_{\theta}[g_{i}(X)]$, based on the known coverages $g_{i}$ provided by the insurance policy. If $X$ follows a compound distribution (i.e. a random sum) then the pure premium does not have a closed form expression but can be approximated to any desired level of precision using simulations.
In the second stage, the loading functions $f_{i}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are approximated using an isotonic regression model, chosen for its ability to maintain the monotonic relationship between pure and commercial premiums—a desirable feature. Additionally, market data is inherently noisy, and isotonic regression provides robustness to outliers, superior to that of simple linear regression, see further discussion on estimation properties of isotonic regression in [15] and references therein. The procedure may be summarized as follows:
@@
The rest of the paper is organized as follows. Section 2 describes the risk model used in this study and discusses insurance pricing principles. Section 3 provides a detailed account of the algorithmic procedure. Our method is presented as an Approximate Bayesian Computation (ABC)-type optimization algorithm, which incorporates a simple isotonic regression model. Section 4 presents the results of a simulation study designed to showcase the performance of our method in a controlled environment. Then, we apply our algorithm on a dataset made of real-world pet insurance rates in Section 5. Finally, we conclude in Section 6 and discuss the perspectives and limits of our methodology for potential applications on other insurance products.
@@ 2.1 Pure premium computation
[…]
where $r\in(0,1]$ is the coverage rate, $d>0$ is the deductible and $l>0$ is the limit. We illustrate the impact of the parameters of the insurance coverage in 1.
@@
Figure 1 shows the pure premiums
@@ 2.2 From pure premiums to commercial premiums
Let $f:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ be a non-decreasing function, such that
@@
The function $f$ is referred to as the loading function. As the commercial premium is a function of the pure premium then we are applying the expectation premium principle. Other premium principles are also possible like the standard deviation principle discussed in Appendix B. A simple loading function is linear in the pure premium as
@@
where $\eta>0$. The loading functions used by insurance companies are unknown to us and vary from one insurance company to the other. We follow up on 1 in 2 where we randomize the linear link between pure and commercial premium.
@@
Take the pure premiums of 1 and apply the following linear loadings
@@
Figure 2 displays the commercial premium as a function of the pure premium.
@@ 3 Market derived insurance ratemaking
The data at hand is a collection of insurance rates $\tilde{p}_{1:n}=\{\tilde{p}_{1},\ldots,\tilde{p}_{n}\}$ associated to insurance coverages $g_{1},\ldots,g_{n}$ for a given risk $X$ within a particular risk class. We suppose that these insurance rates $\mathcal{D}$ were derived according to the model described in 1.
@@
Suppose that the distribution of $X$ is parametrized by an unknown parameter $\theta\in\Theta\subset\mathbb{R}^{d}$. We wish to find the parameters $\theta$ that best explain our data based on a loading function $f$ that attempts to average out the loading functions $f_{1},\ldots,f_{n}$ used by the competitors. Our approach alternates between proposing values for the parameters $\theta$ to calculate the pure premiums and subsequently linking these to the commercial premiums. Section 3.1 discusses a "simpler" optimization problem in which we do not have to handle the loading functions. We use this preliminary problem to introduce our distance function and the notion of model identifiability. Section 3.2 considers the actual problem involving the commercial premiums. The link between pure and commercial premiums is approximated using an isotonic regression model. Regularization terms are included in the distance used to compare observed and model-generated rates to mitigate the identifiability issue. Finally, Section 3.3 refines the accept-reject algorithm laid out in the introduction to search the parameter space in a more efficient way.
@@ 3.1 Optimization problem based on the pure premium
Find $\theta\in\Theta\subset\mathbb{R}^{d}$ to minimize $d\left(p_{1:n},p_{1:n}^{\theta}\right)$, where
@@
$\text{RMSE}\left(p_{1:n},p^{\theta}_{1:n}\right)=\sqrt{\sum_{i=1}^{n}w^{\text{RMSE}}_{i}\left(p_{i}-p_{i}^{\theta}\right)^{2}},$ (4)
@@
$\theta^{\ast}=\underset{\theta\in\Theta}{\operatorname*{arg\,min}}\,\text{RMSE}\left(p_{1:n},p^{\theta}_{1:n}\right).$ (5)
Existence stems from the fact that the parameter space $\Theta$ is compact and the map $\theta\mapsto\text{RMSE}(p_{1:n},p_{1:n}^{\theta})$ is continuous. Uniqueness is more difficult to verify as it depends on the functional $g_{i}$’s. Given the model for $X$ and the insurance coverages, the pure premium does not have an analytical expression making it difficult to show the convexity of (4). A simple necessary condition is that the number of parameters must be smaller than $n$, the number of moments considered. The shape of our insurance coverage function suggests that we may successfully identify the correct parameters, as demonstrated in 3.
@@
We consider the same model as in 1. Recall that the risk has a Poisson-lognormal distribution $X\sim\mathsf{Poisson}(\lambda=3)-\mathsf{LogNorm}(\mu=0,\sigma=1)$ and that $n=100$ insurance coverages are proposed. The rates, deductibles and limit, are set randomly as
@@
Let us assume that $\mu=0$ and compute the pure premium over a grid of values for $\lambda$ and $\sigma$. Figure 3 shows the $\text{RMSE}[p_{1:n},p_{1:n}^{\theta}]$ depending on the value of $\lambda$ and $\sigma$ for $(\lambda,\sigma)\in[0,5]\times[0,2]$.
@@
If uniqueness of the solution cannot be verified, a workarround consists in adding regularization terms to the discrepancy measure. We explore this dircetion in Section 3.2 where the actual problem is treated. The issue is further addressed by uing a particle based optimization methods to search the parameter space and provide a set of admissible candidate parameters. Such an optimization method is presented in Section 3.3. This identifiability issue from an empirical point of view in the online supplementary material [^4] https://github.com/LaGauffre/market_based_insurance_ratemaking/blob/main/latex/supp_material.pdf.
@@ 3.2 Optimization problem based on the commercial premiums
Find $\theta\in\Theta\subset\mathbb{R}^{d}$ and $f:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ to minimize $d\left[\tilde{p}_{1:n},f\left(p_{1:n}^{\theta}\right)\right]$, where the function $f$ is applied elementwise on $p_{1:n}^{\theta}$ and $d(\cdot,\cdot)$ denotes a distance function over the observation space, subject to
@@
$\sum_{i=1}^{n}w_{i}^{\textsc{iso}}(\tilde{p}^{\theta}_{i}-\tilde{p}_{i})^{2}\text{, subject to }\tilde{p}_{i}^{\theta}\leq\tilde{p}^{\theta}_{j}\text{ whenever }p_{i}^{\theta}\leq p_{j}^{\theta},$
@@
    Iterate through the sequence and identify "violations," which occur when the current value is greater than the next value, that is
    $\tilde{p}_{i}^{\ast}>\tilde{p}_{i+1}^{\ast}\text{ for some }i=1,\ldots,n.$
    When a violation is found, adjust the values in the associated segment of the sequence to be the average of the values,
    $\tilde{p}_{i}^{\ast}\leftarrow(\tilde{p}_{i}^{\ast}+\tilde{p}_{i+1}^{\ast})/2,$
    ensuring monotonicity.
@@
We use the isoreg function from $R$ to get the fitted values $\tilde{p}^{\theta}_{i},\text{ }i=1,\ldots,n$. To complete the isotonic regression task we shall find a function $f$ such that $f(p_{i}^{\theta})=\tilde{p}^{\theta}_{i}$. A common choice is a piece-wise constant function that interpolates the $\tilde{p}^{\theta}_{i}$’s. An illustration is provided in 4 and we outline the advantages and limitations associated to this choice in Remark 3.1.
@@
The isotonic fit of the data of 1 and 2 is provided on Figure 4.
@@
When looking at Figure 4, one may object that a simple linear regression model could do the job. This impression is partly due to the (noisy) linear link between pure and commercial premium in 3. Isotonic regression is a non-parametric approach, meaning it doesn’t make strong assumptions about the underlying distribution or functional form of the relationship between variables. This can be advantageous when the true relationship is not well represented by a linear model. We briefly illustrate this fact in Appendix A by looking at the residuals of the linear and isotonic regression and considering a non-linear link function between pure and commercial premium. Furthermore, we believe that ensuring $\tilde{p}_{i}<\tilde{p}_{j}$ when $p_{i}<p_{j}$ is desirable as greater pure premium should be associated to greater commercial premium as a rule of thumb. Isotonic regression aims at satisfying just this condition. The main drawback of isotonic regression is that it can adhere too closely to the data points, risking overfitting. Note also that isotonic regression lacks the interpretability of a simpler, lower-dimensional parametric curve.
@@
$\text{RMSE}\left[\tilde{p}_{1:n},f\left(p^{\theta}_{1:n}\right)\right]=\sqrt{\sum_{i=1}^{n}w^{\text{RMSE}}_{i}\left[\tilde{p}_{i}-f\left(p_{i}^{\theta}\right)\right]^{2}},$ (7)
for candidate risk parameter $\theta$ and the isotonic fit $f$ We seek
$\theta^{\ast}=\underset{\theta\in\Theta}{\operatorname*{arg\,min}}\,\text{RMSE}\left[\tilde{p}_{1:n},f\left(p^{\theta}_{1:n}\right)\right].$
The existence of such $\theta^{\ast}$ is guaranteed because $\theta\mapsto\text{RMSE}\left[\tilde{p}_{1:n},f\left(p^{\theta}_{1:n}\right)\right]$ only takes a finite number of values. Indeed, to each $\theta\in\Theta$ is associated a unique permutation $s^{\theta}\in S_{n}$, where $S_{n}$ denotes the set of all the permutations of $\{1,\ldots,n\}$, such that
@@
leading to a given RMSE value $\text{RMSE}\left[\tilde{p}_{1:n},f\left(p^{\theta}_{1:n}\right)\right]$. Concretely, for $\theta_{1},\theta_{2}\in\Theta$, if it holds that $s^{\theta_{1}}=s^{\theta_{2}}$ then $\text{RMSE}\left[\tilde{p}_{1:n},f(p^{\theta_{1}}_{1:n})\right]=\text{RMSE}\left[\tilde{p}_{1:n},f(p^{\theta_{2}}_{1:n})\right]$. The application $\theta\mapsto s_{n}^{\theta}$ is surjective since $S_{n}^{\Theta}=\{s_{n}^{\theta}\text{ ; }\theta\in\Theta\}$ is finite. The fact that $\Theta$ is a continuous space implies that $\theta^{\ast}$ cannot be unique. The application of isotonic regression offers a straightforward interpretation of our objective, as detailed in Remark 3.2.
@@
At this stage, the optimization problem simplifies to identifying the parameter value $\theta$ corresponding to the most suitable permutation $s^{\theta}$ of the pure premium, which provides the best isotonic fit. We adhere to the guiding principle that a higher pure premium implies a higher commercial premium, as noted in Remark 3.1.
Our problem is an ill-posed inverse problem. Ill-posedness is usually dealt with by adding a regularization to the objective function that one wants to minimize. The ratio of $p/\tilde{p}$ corresponds to what practitioners would call the expected Loss Ratio (LR). Our solution is based on targeting a given loss ratio. The loss ratio is a standard measure to assess the profitability of insurance lines of business. An insurance company that enters a new market is likely to have insights on the loss ratio relative to this market, for example by having informal discussions with reinsurers, brokers or competitors. These feedbacks may translate into the definition of a lower and upper bound denoted by $\text{LR}_{\text{low}}$ and $\text{LR}_{\text{high}}$, respectively. We can then assume that the loss ratios $\text{LR}_{i}=p_{i}/\tilde{p}_{i}$, for $i=1,\ldots,n$, should fall in the range $[\text{LR}_{\text{low}},\text{LR}_{\text{high}}]$, which we refer to as the loss ratio corridor. Assuming that $\text{LR}_{\text{high}}<1$, we ensure both constraints in (6) by adding to our distance (7) two regularization terms defined as
$\text{Reg}_{\text{low}}\left(\tilde{p}_{1:n},p_{1:n}^{\theta}\right)=\sqrt{\sum_{i=1}^{n}w_{i}^{\text{RMSE}}\left(\tilde{p}_{i}-p_{i}^{\theta}\cdot\text{LR}_{\text{low}}^{-1}\right)_{+}^{2}},$
@@
$\text{Reg}_{\text{high}}\left(\tilde{p}_{1:n},p_{1:n}^{\theta}\right)=\sqrt{\sum_{i=1}^{n}w_{i}^{\text{RMSE}}\left(p_{i}^{\theta}\cdot\text{LR}_{\text{high}}^{-1}-\tilde{p}_{i}\right)_{+}^{2}},$
where $(x)_{+}=\max(x,0)$ denotes the positive part of $x$. The distance we consider within 2 is now given by
@@
We illustrate the impact of adding the regularization terms in 5.
@@
We consider the commercial rates of 2. Recall that the commercial premiums are given by
@@
where the pure premiums are those of 3 and
@@
Figure 5 displays the contour plot of the discrepancy between observed and model-generated commercial rates.
@@
When comparing Figure 5a and Figure 5b, we note how beneficial including regularization terms is to identify the true parameter values.
Regularization brings us closer to the scenario described in Section 3.1, where the pure premium is known. This corresponds to the case where $\text{LR}_{\text{low}}=\text{LR}_{\text{high}}=1$. Regularization enables us to exclude large portions of the parameter space associated with nonsensical pure premiums given the commercial premiums. However, careful consideration is required when setting the loss ratio corridor. A narrow loss ratio corridor results in a highly precise estimation, but this estimation may be biased if the loss ratio corridor is misspecified.
The issue of identifiability persists, as multiple parameter values can still solve the optimization problem. This highlights the necessity of using particle-based optimization techniques, as described in Section 3.3, to explore the parameter space. Such techniques return a set of admissible candidate parameters and are less sensitive to initialization. For a more detailed discussion of the identifiability issue, we refer the reader to the online supplementary material [^5] https://github.com/LaGauffre/market_based_insurance_ratemaking/blob/main/latex/supp_material.pdf.

### v2

@@ 1 Introduction
[…]
In the second stage, the loading functions $R+f_{i}$$\mathbb{R}_{+}\to\mathbb{R}_{+}$ are approximated using an isotonic regression model, chosen for its ability to maintain the monotonic relationship between pure and commercial premiums—a desirable feature. Additionally, market data is inherently noisy, and isotonic regression provides robustness to outliers, superior to that of simple linear regression, see further discussion on estimation properties of isotonic regression in [15] and references therein. The procedure may be summarized as follows:
@@
The rest of the paper is organized as follows. Section˜2 describes the risk model used in this study and discusses insurance pricing principles. Section˜3 provides a detailed account of the algorithmic procedure. Our method is presented as an Approximate Bayesian Computation (ABC)-type optimization algorithm, which incorporates a simple isotonic regression model. Section˜4 presents the results of a simulation study designed to showcase the performance of our method in a controlled environment. Then, we apply our algorithm on a dataset made of real-world pet insurance rates in Section˜5. Finally, we conclude in Section˜6 and discuss the perspectives and limits of our methodology for potential applications on other insurance products.
@@ 2.1 Pure premium computation
$g(x)=\min(\max(r\cdot x-d,\,0),\,l),$
where $r\in(0,1]$ is the coverage rate, $d>0$ is the deductible and $l>0$ is the limit. We illustrate the impact of the parameters of the insurance coverage in ˜1.
@@
Figure˜1 shows the pure premiums
@@ 2.2 From pure premiums to commercial premiums
[…]
@@
The function $f$ is referred to as the loading function. As the commercial premium is a function of the pure premium then we are applying the expectation premium principle. Other premium principles are also possible like the standard deviation principle discussed in Appendix˜B. A simple loading function is linear in the pure premium as
@@
where $\eta>0$. The loading functions used by insurance companies are unknown to us and vary from one insurance company to the other. We follow up on ˜1 in ˜2 where we randomize the linear link between pure and commercial premium.
@@
Take the pure premiums of ˜1 and apply the following linear loadings
@@
Figure˜2 displays the commercial premium as a function of the pure premium.
@@ 3 Market derived insurance ratemaking
The data at hand is a collection of insurance rates $\tilde{p}_{1:n}=\{\tilde{p}_{1},\ldots,\tilde{p}_{n}\}$ associated to insurance coverages $g_{1},\ldots,g_{n}$ for a given risk $X$ within a particular risk class. We suppose that these insurance rates $\mathcal{D}$ were derived according to the model described in ˜1.
@@
[…]
@@ 3.1 Optimization problem based on the pure premium
[…]
@@
$\text{RMSE}\left(p_{1:n},p^{\theta}_{1:n}\right)=\sqrt{\sumop\displaylimits_{i=1}^{n}w^{\text{RMSE}}_{i}\left(p_{i}-p_{i}^{\theta}\right)^{2}},$ (4)
@@
[…]
Existence stems from the fact that the parameter space is compact and the map $\theta\mapsto\text{RMSE}(p_{1:n},p_{1:n}^{\theta})$ is continuous. Uniqueness is more difficult to verify as it depends on the functional $g_{i}$’s. Given the model for $X$ and the insurance coverages, the pure premium does not have an analytical expression making it difficult to show the convexity of (4). A simple necessary condition is that the number of parameters must be smaller than $n$, the number of moments considered. The shape of our insurance coverage function suggests that we may successfully identify the correct parameters, as demonstrated in ˜3.
@@
We consider the same model as in ˜1. Recall that the risk has a Poisson-lognormal distribution $X\sim\mathsf{Poisson}(\lambda=3)-\mathsf{LogNorm}(\mu=0,\sigma=1)$ and that $n=100$ insurance coverages are proposed. The rates, deductibles and limit, are set randomly as
@@
Let us assume that $\mu=0$ and compute the pure premium over a grid of values for $\lambda$ and $\sigma$. Figure˜3 shows the $\text{RMSE}[p_{1:n},p_{1:n}^{\theta}]$ depending on the value of $\lambda$ and $\sigma$ for $(\lambda,\sigma)\in[0,5]\times[0,2]$.
@@
If uniqueness of the solution cannot be verified, a workarround consists in adding regularization terms to the discrepancy measure. We explore this dircetion in Section˜3.2 where the actual problem is treated. The issue is further addressed by uing a particle based optimization methods to search the parameter space and provide a set of admissible candidate parameters. Such an optimization method is presented in Section˜3.3. This identifiability issue from an empirical point of view in the online supplementary material [^4] https://github.com/LaGauffre/market_based_insurance_ratemaking/blob/main/latex/supp_material.pdf.
@@ 3.2 Optimization problem based on the commercial premiums
[…]
@@
$\sumop\displaylimits_{i=1}^{n}w_{i}^{\textsc{iso}}(\tilde{p}^{\theta}_{i}-\tilde{p}_{i})^{2}\text{, subject to }\tilde{p}_{i}^{\theta}\leq\tilde{p}^{\theta}_{j}\text{ whenever }p_{i}^{\theta}\leq p_{j}^{\theta},$
@@
    Iterate through the sequence and identify "violations," which occur when the current value is greater than the next value, that is ~p_i^∗>~p_i+1^∗ for some i=1,…, n. When a violation is found, adjust the values in the associated segment of the sequence to be the average of the values, ~p_i^∗←(~p_i^∗+ ~p_i+1^∗)/2, ensuring monotonicity.
@@
We use the isoreg function from $R$ to get the fitted values $\tilde{p}^{\theta}_{i},\text{ }i=1,\ldots,n$. To complete the isotonic regression task we shall find a function $f$ such that $f(p_{i}^{\theta})=\tilde{p}^{\theta}_{i}$. A common choice is a piece-wise constant function that interpolates the $\tilde{p}^{\theta}_{i}$’s. An illustration is provided in ˜4 and we outline the advantages and limitations associated to this choice in Remark˜3.1.
@@
The isotonic fit of the data of ˜1 and ˜2 is provided on Figure˜4.
@@
When looking at Figure˜4, one may object that a simple linear regression model could do the job. This impression is partly due to the (noisy) linear link between pure and commercial premium in ˜3. Isotonic regression is a non-parametric approach, meaning it doesn’t make strong assumptions about the underlying distribution or functional form of the relationship between variables. This can be advantageous when the true relationship is not well represented by a linear model. We briefly illustrate this fact in Appendix˜A by looking at the residuals of the linear and isotonic regression and considering a non-linear link function between pure and commercial premium. Furthermore, we believe that ensuring $\tilde{p}_{i}<\tilde{p}_{j}$ when $p_{i}<p_{j}$ is desirable as greater pure premium should be associated to greater commercial premium as a rule of thumb. Isotonic regression aims at satisfying just this condition. The main drawback of isotonic regression is that it can adhere too closely to the data points, risking overfitting. Note also that isotonic regression lacks the interpretability of a simpler, lower-dimensional parametric curve.
@@
$\text{RMSE}\left[\tilde{p}_{1:n},f\left(p^{\theta}_{1:n}\right)\right]=\sqrt{\sumop\displaylimits_{i=1}^{n}w^{\text{RMSE}}_{i}\left[\tilde{p}_{i}-f\left(p_{i}^{\theta}\right)\right]^{2}},$ (7)
for candidate risk parameter $\theta$ and the isotonic fit $f$ . We seek
[…]
@@
[…]
@@
At this stage, the optimization problem simplifies to identifying the parameter value $\theta$ corresponding to the most suitable permutation $s^{\theta}$ of the pure premium, which provides the best isotonic fit. We adhere to the guiding principle that a higher pure premium implies a higher commercial premium, as noted in Remark˜3.1.
Our problem is an ill-posed inverse problem. Ill-posedness is usually dealt with by adding a regularization to the objective function that one wants to minimize. The ratio of $p/\tilde{p}$ corresponds to what practitioners would call the expected Loss Ratio (LR). Our solution is based on targeting a given loss ratio. The loss ratio is a standard measure to assess the profitability of insurance lines of business. An insurance company that enters a new market is likely to have insights on the loss ratio relative to this market, for example by having informal discussions with reinsurers, brokers or competitors. These feedbacks may translate into the definition of a lower and upper bound denoted by $\text{LR}_{\text{low}}$ and $\text{LR}_{\text{high}}$, respectively. We can then assume that the loss ratios $\text{LR}_{i}=p_{i}/\tilde{p}_{i}$, for $i=1,\ldots,n$, should fall in the range $\text{LR}_{\text{low}},\text{LR}_{\text{high}}]$, which we refer to as the loss ratio corridor. Assuming that $\text{LR}_{\text{high}}<1$, we ensure both constraints in (6) by adding to our distance (7) two regularization terms defined as
$\text{Reg}_{\text{low}}\left(\tilde{p}_{1:n},p_{1:n}^{\theta}\right)=\sqrt{\sumop\displaylimits_{i=1}^{n}w_{i}^{\text{RMSE}}\left(\tilde{p}_{i}-p_{i}^{\theta}\cdot\text{LR}_{\text{low}}^{-1}\right)_{+}^{2}},$
@@
$\text{Reg}_{\text{high}}\left(\tilde{p}_{1:n},p_{1:n}^{\theta}\right)=\sqrt{\sumop\displaylimits_{i=1}^{n}w_{i}^{\text{RMSE}}\left(p_{i}^{\theta}\cdot\text{LR}_{\text{high}}^{-1}-\tilde{p}_{i}\right)_{+}^{2}},$
where $(x)_{+}=\max(x,0)$ denotes the positive part of $x$. The distance we consider within ˜2 is now given by
@@
We illustrate the impact of adding the regularization terms in ˜5.
@@
We consider the commercial rates of ˜2. Recall that the commercial premiums are given by
@@
where the pure premiums are those of ˜3 and
@@
Figure˜5 displays the contour plot of the discrepancy between observed and model-generated commercial rates.
@@
When comparing Figure˜5a and Figure˜5b, we note how beneficial including regularization terms is to identify the true parameter values.
Regularization brings us closer to the scenario described in Section˜3.1, where the pure premium is known. This corresponds to the case where $\text{LR}_{\text{low}}=\text{LR}_{\text{high}}=1$. Regularization enables us to exclude large portions of the parameter space associated with nonsensical pure premiums given the commercial premiums. However, careful consideration is required when setting the loss ratio corridor. A narrow loss ratio corridor results in a highly precise estimation, but this estimation may be biased if the loss ratio corridor is misspecified.
The issue of identifiability persists, as multiple parameter values can still solve the optimization problem. This highlights the necessity of using particle-based optimization techniques, as described in Section˜3.3, to explore the parameter space. Such techniques return a set of admissible candidate parameters and are less sensitive to initialization. For a more detailed discussion of the identifiability issue, we refer the reader to the online supplementary material [^5] https://github.com/LaGauffre/market_based_insurance_ratemaking/blob/main/latex/supp_material.pdf.